# Disentangling spatio-temporal knowledge for weakly supervised object detection and segmentation in surgical video

### v1

@@ Abstract
Weakly supervised video object segmentation (WSVOS) enables the identification of segmentation maps without requiring an extensive training dataset of object masks, relying instead on coarse video labels indicating object presence. Current state-of-the-art methods either require multiple independent stages of processing that employ motion cues or, in the case of end-to-end trainable networks, lack in segmentation accuracy, in part due to the difficulty of learning segmentation maps from videos with transient object presence. This limits the application of WSVOS for semantic annotation of surgical videos where multiple surgical tools frequently move in and out of the field of view, a problem that is more difficult than typically encountered in WSVOS. This paper introduces Video Spatio-Temporal Disentanglement Networks (VDST-Net), a framework to disentangle spatiotemporal information using semi-decoupled knowledge distillation to predict high-quality class activation maps (CAMs). A teacher network designed to resolve temporal conflicts when specifics about object location and timing in the video are not provided works with a student network that integrates information over time by leveraging temporal dependencies. We demonstrate the efficacy of our framework on a public reference dataset and on a more challenging surgical video dataset where objects are, on average, present in less than 60% of annotated frames. Our method outperforms state-of-the-art techniques and generates superior segmentation masks under video-level weak supervision.
@@ 1 Introduction
In Weakly Supervised Video Object Segmentation (WSVOS), object labels are only available at video level, introducing uncertainty about object presence in each frame. Some approaches to WSVOS leverage motion cues  [40, 51]. However, these methods often lack robustness against background motion and stationary objects. A recent end-to-end framework [1] has achieved state-of-the-art performance in object detection on two challenging datasets. In these datasets, the primary objects or concepts present in the video are labeled and are constantly present in most frames. We refer to this type of problem as Consistent Object Presence (COP), in contrast to Transient Object Presence (TOP) where no assumptions are made about object constancy (Figure 1). The latter case covers the significant variation in temporal presence of objects in surgical videos, where annotations are often based on the start and end times of surgical steps with annotated tools intermittently moving in and out of view. In Transient Object Presence (WSVOS-TOP), the algorithm must thus learn only from object classes without any implicit or explicit information on location, size, or temporal presence, and without assumptions about a specific count of object classes per frame.
@@
In this work, we propose a method for WSVOS that addresses challenges with both COP and TOP problems. Our contributions include:
@@
    Experiments on YouTube-Objects (v2.2) demonstrate that our method outperforms state-of-the-art approaches in classic WSVOS-COP.
@@ 2.2 Weakly supervised detection and segmentation for video
Weakly supervised video segmentation can be categorized into three types of tasks, as shown in Figure 1. In the first type, frame-level object presence labels are provided. Research in this area often incorporates motion constraints to enhance prediction accuracy, as demonstrated in studies such as [30, 27, 50]. [32] employs LSTM to reason across time for detection and tracking in surgical videos.[13] adapts a 3D CNN to incorporate temporal dependencies for detection in videos. The second type of WSVOS has object labels available only at the video-level, thus not explicitly determining object presence in a particular frame. However, an implicit assumption in the literature is that of a Consistent Object Presence (WSVOS-COP) setting, namely that the labeled objects are present throughout the video duration, treating occasional frames with missing objects as noise. This situation is exemplified by well-known reference datasets such as the YouTube-Objects datasets [17], which in addition contain only one object class per video. Consequently, this setup is close to a Type I problem as frame content can be inferred from video-level annotations. Recent work on TCAM [1] extracts the most reliable pseudo-semantic supervision signal in video to train a subsequent network for detection in individual images, efficiently addressing object detection for this case, while still showing limited robustness to smaller objects and temporal inconsistencies. In the third type (Type III), labels merely indicate the general presence of objects throughout an entire video, despite each object possibly appearing only temporarily. This is depicted in the right column of Figure 1. Consequently, the presence label assigned to a video may not accurately reflect a large subset of frames, introducing a significant challenge in learning generalized representations due to the increased uncertainty. Using motion cues could not resolve such an issue when including scenarios where the camera can occasionally move while the objects stay still. To the best of our knowledge, our work is the first to address this type of TOP data and train an end-to-end network to predict segmentation masks for the video as a whole.
@@ 2.3 Network reasoning across time and knowledge distillation
3D CNN [41], recurrent neural networks (RNN) [36], and temporal Transformers [47] are widely used architectures that capture time dependency and have been applied to various video and volumetric imaging tasks [28, 54, 31]. However, due to the lack of direct supervision signals, learning spatiotemporal features for weakly supervised segmentation is challenging. Knowledge Distillation (KD) [10] using a teacher-student network was originally designed for model compression and can gradually improve network learning capacity [6]. This property allows KD to enhance learning in weakly or semi-supervised tasks. L2G [14] applied KD to extend CAM from local to global. SCD [46] uses KD to incorporate feature correspondence for CAM refinement in images. Seco [48] employs KD to decouple occurrences for refining CAM in images. Our method employs KD to address the challenges posed by ill-posed video-level weak supervision, which causes activation conflicts leading to learning of features that are not associated with the correct class. We use a constrained teacher to allow the network not only to produce correct activation, but also to refine the backbone to learn generalized features for new domain-specific data. Decoupled KD [52] has proven to be a versatile method for flexibly distilling meaningful knowledge. Our design, which could be categorized as being based on semi-decoupled KD, is tailored for video weakly supervised segmentation. Rather than relying on soft labels from the teacher, we enable the teacher’s predictions while utilizing ground truth to form gates for student supervision. This approach allows the student to learn effectively from an imperfect teacher while still gaining temporal reasoning abilities.
Figure 2: Our VDST-Net approach deploying knowledge distillation to disentangle spatial and temporal information for weakly supervised learning with video-level labels.
@@ 3 Methods
Figure 2 illustrates the proposed VDST-Net framework. The core of the framework consists of a teacher-student network pair designed to disentangle spatial and temporal knowledge. Both modules utilize a spatio-temporal class activation map (ST-CAM) and share input from a ViT [4]. The primary distinction between the teacher and student lies in their upper-layer feature extraction: the teacher employs a Multilayer Perceptron (MLP) module, while the student module uses a temporal convolutional network. This design constrains the teacher’s temporal interaction to overcome activation conflicts, while gradually increasing student’s learning capability to reason over time. We utilize a semi-decoupled distillation mechanism, supervising teacher and student models with identical video labels while enabling efficient knowledge transfer via gated activation maps. Despite the teacher’s powerful backbone and its ability to generate activation maps with a high degree of accuracy it still has detection gaps. Although the resulting supervision signal is not perfect, semi-decoupled distillation allows the student to use the teacher CAMs information efficiently while adding temporal context.
@@ 3.1 Video encoder
A ViT encoder is used to encode patches from individual frames of the input video $V\in{\mathbb{R}}^{W\times H\times D}$, producing a stack of token embeddings reshaped as $T\in{\mathbb{R}}^{w\times h\times D\times C_{e}}$, where $w=W/P$, $h=H/P$, and $P$ denotes the width/height of a patch, and $C_{e}$ is the embedding dimension. $T$ captures features with strict temporal correspondence (maintaining the temporal dimension $D$) and high spatial correspondence.
@@ 3.2 Teacher module
$T$ is further processed by a MLP block that linearly projects $T$ to $C$ channels of a feature map $F^{t}\in{\mathbb{R}}^{w\times h\times D\times C}$, maintaining spatio-temporal correspondence. Following a line of work on ViT-based 2D CAM [7, 45], spatial pooling is applied to $F^{t}$ to obtain a slice-wise feature map $F_{s}^{t}\in{\mathbb{R}}^{1\times 1\times D\times C}$. Then temporal pooling is applied to acquire a final feature map for video classification $F_{v}^{t}\in{\mathbb{R}}^{1\times 1\times 1\times C}$. Inspired by weakly supervised temporal scene localization [38], a ranked top-$k$ average pooling is adopted. This approach first ranks all the spatial patch features by magnitude, and then selects the top-$k_{1}$ patches for average pooling. Then ranked top-$k_{2}$ is applied to select the temporal feature. It balances between max pooling and average pooling; a fine-tuned optimal k-value could be correlated to object-to-background ratio and within-video object presence frequency. Finally, with a fully connected classification layer $\mathcal{F_{C}}^{t}$, the final video prediction $P_{v}^{t}\in{\mathbb{R}}^{1\times 1\times 1\times N}$ is obtained by $P_{v}^{t}=\mathcal{F_{C}}^{t}(F_{v}^{t})$, where $N$ is the number of classes. Plugging $\mathcal{F_{C}}^{t}$ channel-wise to feature map $F^{t}$ results in a multi-channel ST-CAM $M^{t}\in{\mathbb{R}}^{w\times h\times D\times N}$.
@@ 3.3 Student module
Employing an MLP for encoding of embeddings $T$ processes each frame in isolation, thus eliminating temporal conflicts. However, it fails to exploit temporal coherence to enhance prediction robustness. Temporal convolutional layers on the other hand can infer associations between frames, yet will introduce errors due to the absence of frame-level annotation. To mitigate this issue, we utilize a semi-decoupled KD technique that refines a student network with higher capacity through efficiently selected supervision signals. In this context, we designate a temporal convolution-based network as the student module. The student network adopts the same spatio-temporal pooling mechanism from the teacher and trains an additional fully connected classification layer $\mathcal{F_{C}}^{s}$. This layer produces the student’s video prediction $P_{v}^{s}\in{\mathbb{R}}^{1\times 1\times 1\times N}$ and the student ST-CAM $M^{s}\in{\mathbb{R}}^{w\times h\times D\times N}$.
3.4 Semi-decoupled distillation and training loss
To help the backbone learn generalized features at different scales, features from each MLP layer in the teacher module are randomly downsampled during training with a probability of 0.5. Next, we employ a semi-decoupled [52] approach to knowledge distillation (KD), where we use video ground truth labels to supervise both teacher and student networks’ predictions while applying KD on activation maps. The teacher module is trained solely on a multi-label soft margin classification loss $\mathcal{L}_{ml}(P_{v}^{t},P_{g})$ using ground truth video labels $P_{g}\in{\mathbb{R}}^{1\times N}$. For the student module, in addition to the classification loss, we include a KD loss to further refine the student’s activation map. This loss is computed by considering only the true positive channel (TPC) and positive slice predictions from the teacher network.
Initially, a teacher slice prediction $P_{s}^{t}\in{\mathbb{R}}^{1\times 1\times D\times N}$ is obtained by applying maximum spatial pooling to the teacher ST-CAM. Subsequently, we create a TPC slice kernel $K_{tp}\in{\mathbb{R}}^{1\times 1\times D\times N}$ through the operation $K_{tp}=P_{s}^{t}\otimes r(P_{g})$, where $r(\cdot)$ represents a repeating operation to match dimensions for element-wise multiplication $\otimes$. A final gated KD loss is calculated by:
${l}_{KD}=\mathcal{L}_{mse}(M^{s}\otimes r(K_{tp}),M^{t}\otimes r(K_{tp}))\enspace,$ (1)
where $\mathcal{L}_{mse}$ is a Mean Squared Error loss function, and $r(\cdot)$ here matches the dimension of $K_{tp}$ to $M^{t}$ and $M^{s}$. In analogy, this operation resembles focusing exclusively on frames where true positive pseudo masks are predicted, akin to singling out specific instances for loss calculation. The final loss used for optimizing the student module is given by:
${l}_{s}=\mathcal{L}_{ml}(P_{v}^{s},P_{g})+\alpha{l}_{KD}\enspace.$ (2)
@@ 4 Experiments
Table 1: Detection performance on YouTube-Objects [33] dataset under CorLoc Metrics.

### v3

@@ Abstract
Weakly supervised video object segmentation (WSVOS) enables the identification of segmentation maps without requiring extensive annotations of object masks, relying instead on coarse video labels indicating object presence. Weakly supervised semantic segmentation of objects in surgical videos is however more challenging due to a complex interaction of multiple transient objects, such as surgical tools moving in and out of the surgical field. In this scenario, state-of-the-art WSVOS methods struggle to learn accurate segmentation maps. We address this problem by introducing Video Spatio-Temporal Disentanglement Networks (VDST-Net), a framework to disentangle complex spatiotemporal object interactions using semi-decoupled knowledge distillation to predict high-quality class activation maps (CAMs). A teacher network is designed to help a temporal-reasoning student network resolve activation conflicts, as the student leverages temporal dependencies when specifics about object location and timing in the video are not provided. We demonstrate the efficacy of our framework on a challenging surgical video dataset where objects are, on average, present in less than 60% of annotated frames, and compare our method to state-of-the-art methods on surgical data and on a public dataset commonly used to benchmark WSVOS. Our method outperforms state-of-the-art techniques and generates accurate segmentation masks under video-level weak supervision.
@@ 1 Introduction
In the more difficult problem of Weakly Supervised Video Object Segmentation (WSVOS), object presence labels are only available at video level, while no image level presence or segmentation supervision signal is provided. This introduces higher uncertainty about object presence in each frame. Some approaches to WSVOS leverage motion cues  [42, 56]. However, these methods often lack robustness against background motion and stationary objects. A recent end-to-end framework [1] has achieved state-of-the-art performance in object detection on two challenging datasets. In these datasets, the primary objects or concepts present in the video are labeled and are constantly present in most frames. We refer to this type of problem as Consistent Object Presence (COP), in contrast to Transient Object Presence (TOP) where no assumptions are made about object constancy (Figure 1). The latter case covers the significant variation in temporal presence of objects in surgical videos, where annotations are often based on the start and end times of surgical steps with annotated tools intermittently moving in and out of view. In Transient Object Presence (WSVOS-TOP), the algorithm must thus learn only from object classes without any implicit or explicit information on location, size, or temporal presence, and without assumptions about a specific count of object classes per frame.
@@
In order to address these challenges, we propose a novel method for WSVOS that excels at the TOP problem on surgical data while also achieving state-of-the-art performance in COP scenarios. Our contributions include:
@@
    Experiments on YouTube-Objects (v2.2) demonstrate that our method also outperforms state-of-the-art approaches in classic WSVOS-COP.
@@ 2.2 Weakly supervised detection and segmentation for video
WSVOS can be categorized into three types of tasks, as shown in Figure 1. In the first type (Type I), frame-level object presence labels are provided. Research in this area often incorporates motion constraints to enhance prediction accuracy, as demonstrated in studies such as [32, 29, 53]. [34] employs LSTM to reason across time for detection and tracking in surgical videos.[14] adapts a 3D CNN to incorporate temporal dependencies for detection in videos. The second type of WSVOS (Type II) has object labels available only at the video-level, thus not explicitly determining object presence in a particular frame. However, a common assumption in the literature in the WSVOS-COP setting is that labeled objects are present throughout the entire video, with frames missing these objects treated as noise. This scenario is exemplified by well-known datasets such as YouTube-Objects[18].Consequently, this setup is close to a Type I problem as frame content can be inferred from video-level annotations. Recent work on TCAM [1] extracts the most reliable pseudo-semantic supervision signal in video to train a subsequent network for detection in individual images, efficiently addressing object detection for this case, while still showing limited robustness to smaller objects and temporal inconsistencies. In the third type (Type III), labels merely indicate the general presence of objects throughout an entire video, despite each object possibly appearing only temporarily. This is depicted in the right column of Figure 1. Consequently, the presence label assigned to a video may not accurately reflect a large subset of frames, introducing higher uncertainty that leads to challenge in learning generalized representations. Using motion cues could not resolve such an issue when including scenarios where the camera can occasionally move while the objects stay still. Existing semantic segmentation methods that use sparse or coarse labels to segment surgical video only focus on the WSSS problem with presence labels for individual frames available[34, 51, 27]. Zero-shot approaches rely on SAM and a prompting strategy[54]. However, no attempts have been made to apply WSVOS to surgical videos or similar TOP scenarios using a single class presence label per video. To the best of our knowledge, our work is the first to address this type of TOP data and train an end-to-end network to predict segmentation masks for the video as a whole.
@@ 2.3 Network reasoning across time and knowledge distillation
3D CNN [43], recurrent neural networks (RNN) [38], and temporal Transformers [49] are widely used architectures that capture time dependency and have been applied to various video and volumetric imaging tasks [30, 59, 33]. However, due to the lack of direct supervision signals, learning spatiotemporal features for weakly supervised segmentation is difficult. Knowledge Distillation (KD) [11] using a teacher-student network was originally designed for model compression and can gradually improve network learning capacity [6]. This property allows KD to enhance learning in weakly or semi-supervised tasks. L2G [15] applied KD to extend CAM from local to global. SCD [48] uses KD to incorporate feature correspondence for CAM refinement in images. Seco [50] employs KD to decouple occurrences for refining CAM in images. Our method employs KD to address the challenges of ill-posed video-level weak supervision, which causes activation conflicts of features that are not associated with the correct class.
We use a constrained teacher to allow the network not only to produce correct activation, but also to refine the backbone to learn generalized features for new domain-specific data. Decoupled KD [57] has proven to be a versatile method for flexibly distilling meaningful knowledge. Our design, which could be categorized as being based on semi-decoupled KD, is tailored for video weakly supervised segmentation. Rather than relying on soft labels from the teacher, we enable the teacher’s predictions while utilizing ground truth to form gates for student supervision. This approach allows the student to learn effectively from an imperfect teacher while still gaining temporal reasoning abilities.
Figure 2: Left: Our VDST-Net approach deploying knowledge distillation to disentangle spatial and temporal information for weakly supervised learning with video-level labels. Ground truth video presence labels $P_{g}$ provide supervision to both teacher and student, while knowledge in activation map $M^{t}$ is transferred from teacher to student. Right: Ranked spatial and temporal pooling captures information about multiple objects in a frame while filtering out spurious information from frames where a target object is missing.
@@ 3 Methods
Figure 2 (left) illustrates the proposed VDST-Net framework. The core of the framework consists of a teacher-student network pair designed to disentangle spatial and temporal knowledge. Both modules utilize a spatio-temporal class activation map (ST-CAM) and share input from ViT [4] encoded frames and video-level class labels as the supervision signal. The primary distinction between the teacher and the student lies in their upper-layer feature extraction modules: the teacher employs a Multilayer Perceptron (MLP), while the student employs a temporal convolutional network (TCN). This design intentionally restricts the teacher from temporal interactions, reducing the potential for activation conflicts between frames, while progressively increasing the student’s ability to reason across time. We utilize a semi-decoupled online distillation approach, where both teacher and student modules are trained using the same video labels, and efficient knowledge transfer is facilitated through gated activation maps. While the teacher module has a powerful backbone capable of generating highly accurate activation maps, it still exhibits detection gaps and cannot consistently provide fully reliable knowledge for distillation. To resolve these limitations, the semi-decoupled distillation allows the student module to selectively learn from the teacher’s gated outputs under the constraints of video-level ground truth, while enhancing temporal consistency.
@@ 3.1 Video encoder
A ViT encoder is used to encode patches from individual frames of the input video $V\in{\mathbb{R}}^{W\times H\times D}$, producing a stack of token embeddings reshaped as $X\in{\mathbb{R}}^{w\times h\times D\times C_{e}}$, where $w=W/P$, $h=H/P$, and $P$ denotes the width/height of a patch, and $C_{e}$ is the embedding dimension. $X$ captures features with strict temporal correspondence (maintaining the temporal dimension $D$) and high spatial correspondence.
@@ 3.2 Teacher module
As depicted in the blue block of Fig 2,video embedding tensor $X$ is further processed by a MLP block that linearly projects $X$ a feature map $F^{t}\in{\mathbb{R}}^{w\times h\times D\times C}$ with $C$ channels , maintaining spatio-temporal correspondence. Following a line of work on ViT-based 2D CAM [7, 47], spatial pooling is applied to $F^{t}$ to obtain a slice-wise feature map $F_{s}^{t}\in{\mathbb{R}}^{1\times 1\times D\times C}$. Then, temporal pooling is applied to acquire a final feature map for video classification $F_{v}^{t}\in{\mathbb{R}}^{1\times 1\times 1\times C}$. Inspired by weakly supervised temporal scene anomaly localization [40] we use a ranked top-$k$ average pooling:
${{{{F_{s}^{t,i}=\frac{1}{k_{1}}\sum\nolimits_{j=1}^{k_{1}}{F^{t,r_{j}^{i}}}\enspace}}}}$ (1)
where $i\in[1,D]$ is the temporal index of feature $F_{s}^{t}$. $r_{j}^{i}\in[1,H\times W]$ is the ranked spatial index of $F^{t}$ at a given temporal slice. This approach first ranks all the spatial patch features by magnitude, and then selects the top-$k_{1}$ patches for average pooling. Following a similar strategy, ranked temporal pooling is applied as:
${{{{F_{v}^{t}=\frac{1}{k_{2}}\sum\nolimits_{l=1}^{k_{2}}{F_{s}^{t,r_{l}}}\enspace}}}}$ (2)
where $r_{l}\in[1,D]$ is the ranked temporal index of $F_{s}^{t}$. $k_{2}$ is used to select top temporal features. The ranked spatio-temporal pooling process is illustrated in the Fig. 2 (right). Maximum pooling or global average pooling can be seen as extreme cases of ranked pooling (i.e. $k_{2}$=1 or $k_{2}=D$), which encourage activating either on the most salient feature location or whole video, causing low accuracy in the surgical video TOP scenario. Ranked pooling design can mitigate such issues and force the ST-CAM to activate on the proper temporal slices and area of interest. Finally, with a fully connected classification layer $\mathcal{F_{C}}^{t}$, the final video prediction $P_{v}^{t}\in{\mathbb{R}}^{1\times 1\times 1\times N}$ is obtained by $P_{v}^{t}=\mathcal{F_{C}}^{t}(F_{v}^{t})$, where $N$ is the number of classes. Plugging $\mathcal{F_{C}}^{t}$ channel-wise to feature map $F^{t}$ results in a multi-channel ST-CAM $M^{t}\in{\mathbb{R}}^{w\times h\times D\times N}$. To help the backbone learn generalized features at different scales, after each MLP layer, features are randomly downsampled during training with a probability of 0.5.
@@ 3.3 Student module
Employing an MLP for encoding of embeddings $X$, the teacher module processes each frame in isolation, thus eliminating temporal conflicts. However, this fails to exploit the temporal coherence to enhance prediction robustness. We thus designate a temporal convolution-based network in the student module. As depicted in Fig. 2, the student network adopts the same spatio-temporal pooling mechanism from the teacher and trains an additional fully connected classification layer $\mathcal{F_{C}}^{s}$. This layer produces the student’s video prediction $P_{v}^{s}\in{\mathbb{R}}^{1\times 1\times 1\times N}$ and the student ST-CAM $M^{s}\in{\mathbb{R}}^{w\times h\times D\times N}$. The student’s TCN can infer associations between frames, yet will introduce errors due to the absence of frame-level annotation. To mitigate this issue, we utilize a semi-decoupled KD technique that refines a student network with higher capacity through efficiently selected supervision signals.
3.4 Training loss and semi-decoupled knowledge distillation
The teacher module optimizes on a multi-label soft margin classification loss $\mathcal{L}_{ml}(P_{v}^{t},P_{g})$ utilizing ground truth video labels $P_{g}\in\mathbb{R}^{1\times N}$. The student module optimizes the same classification loss and a KD loss that further refines the student’s activation map. We employ a semi-decoupled [57] approach to knowledge distillation where weak video ground truth labels supervise both teacher and student networks’ predictions while applying KD through activation maps. KD was used to transfer teacher’s prediction logits as soft labels[11, 6], or attention[55, 8] ; our approach is closer to latter, as it uses attention maps to guide the student[8]. KD loss is computed by considering only the true positive channel (TPC) and positive slice predictions from the teacher network.
Initially, a teacher slice prediction $P_{s}^{t}\in{\mathbb{R}}^{1\times 1\times D\times N}$ is obtained by applying maximum spatial pooling to the teacher ST-CAM. Subsequently, we create a TPC slice kernel $K_{tp}\in{\mathbb{R}}^{1\times 1\times D\times N}$ through the operation $K_{tp}=P_{s}^{t}\otimes r(P_{g})$, where $r(\cdot)$ represents a broadcast operation to match dimensions for element-wise multiplication $\otimes$. A final gated KD loss is calculated by:
${l}_{KD}=\mathcal{L}_{mse}(M^{s}\otimes r(K_{tp}),M^{t}\otimes r(K_{tp}))\enspace$ (3)
where $\mathcal{L}_{mse}$ is a Mean Squared Error loss function, and $r(\cdot)$ here matches the dimension of $K_{tp}$ to $M^{t}$ and $M^{s}$. In analogy, this operation focusing exclusively on frames where true positive pseudo masks are predicted, akin to singling out specific instances for loss calculation. The final loss used for optimizing the student module is given by:
${l}_{s}=\mathcal{L}_{ml}(P_{v}^{s},P_{g})+\alpha{l}_{KD}\enspace$ (4)
@@ 4 Experiments
4.1 Datasets
To investigate the characteristics of our proposed method, we utilize two distinct video dataset collections: Cholec80[45] dataset and YouTube-Objects video dataset (v2.2) [18]. The Cholec80 dataset consists of 80 cholecystectomy surgery videos, each downsampled to 1 FPS, with presence labels of seven surgical tools ($N=7$) in each frame. We create a variant TOP video dataset based on Cholec80. First these videos are split into 30-second clips without overlap, merging the tool labels across frames to generate a video-level tool presence binary vector label for each clip. The test set is based on CholecSeg8K [12] which contains 8K images from 17 videos in Cholec80 with segmentation labels for instruments and anatomy. From this dataset we derive 100 video clips containing 3K frames. To prevent contamination between testing and training data, 5,296 training clips are sampled from from 63 Cholec80 videos not included in CholecSeg8k. We calculate the tool frequencies across clips using frame per clip (FPC) which denotes the percentage of frames in a clip with the tool present. We also test our method on YouTube-objects to demonstrate the generalizability and robustness of our proposed method across different types of videos.YouTube-objects dataset is a well-known benchmark in the computer vision community which contains 106 videos for training, and 49 videos for testing, where ten categories ($N=10$) of objects are covered. Each video is split into short-duration clips, and in each clip a few frames are annotated with a bounding box for evaluation. YouTube-Objects dataset has nearly 100% FPC for each clip and thus represents the second case (Type II – WSVOS-COP) demonstrated in Figure 1. Most categories in Cholec80 based dataset appear on average in less than 60% of the frames of a particular video labeled with that category, which makes it a much more challenging dataset that represents a WSVOS-TOP (Type III) problem.
4.2 Experimental Setup and Evaluation Metrics
4.3 Surgical video tool segmentation performance
Table 1: Comparison to state-of-the-art weakly supervised methods with either frame-level supervision or video-level supervision. Video (V) and Frame (F) accuracy (AC).
MCTformer [47] 98.27 97.63 43.14 50.07
LayerCAM[16] 95.67 94.87 31.97 75.59
XGradCAM[5] 92.21 90.21 32.13 53.53
TCAM [1] 97.08 95.10 28.22 54.60
T only (ours) 98.26 97.56 47.50 48.72
T-S fusion (ours) 97.11 97.23 48.64 39.61
VDST-Net (ours) 96.82 98.23 61.80 28.10
On surgical videos, we compare our method to frame-level weakly supervised methods MCTformer [47], LayerCAM [16], and XGradCAM [5], all trained with frame-level labels. Models were adapted to our dataset using their open-source code implementations. We also compare our method to the video-level supervised method TCAM [1], where both our method and TCAM are trained with more challenging video-level class labels. We test our model using teacher (T) only, teacher-student (T-S) fusion, and full VDST-Net variations. For the surgical video data we utilize pre-trained Masked Auto-Encoder (MAE) ViT backbone [10] which is tuned on a large scale data for segmentation[19] before it is adapted to the new domain (Cholec80) data.
4.4 Analysis on Youtube video object detection Perfomance
Table 2: Detection performance on YouTube-Objects [35] dataset under CorLoc Metrics. S=Student, T=Teacher